Calcium hydroxide is reacted with hydrochloric acid. What is the weight of calcium chloride produced if 0.15N of hydrochloric acid is used and 26 kilograms of calcium hydroxide are the raw materials?
1
Expert's answer
2020-12-06T10:35:38-0500

Ca(OH)2 + 2HCl —> CaCl2 + 2H2O


0.15N (HCl) = 0.015kg = 15g

No of moles of HCl = 15/(2*35.5) = 15/71 = 0.211mol/dm3


With HCl as the limiting reagent

and molar mass of CaCl2 as 111g/mol


Mass of CaCl2 therefore formed = 0.211*111 = 23.45g = 0.023kg


The weight of CaCl2 is 0.23N